OpenStudy (anonymous):

Emily wants to hang a painting in a gallery but the painting and frame must have an area of 31 square feet. The painting is 5 feet wide by 6 feet long. The frame must have an equal width of x on each side. Which quadratic equation represents the area of the painting and frame combined?

OpenStudy (anonymous):

4x2 + 22x − 1 = 0 4x2 + 22x + 31 = 0 x2 + 11x − 1 = 0 x2 + 11x + 31 = 0

OpenStudy (anonymous):

then the width would be 5+2x and length would be 6+2x multiple those together to get the area and the area has to be equal to 31 ft^2

OpenStudy (anonymous):

do you understand where i got the length and width?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

a little